### v2

# Embedding ample semigroups as $(2,1,1)$-subalgebras of inverse semigroups

Abstract.
The problem of embedding an ample semigroup in an inverse semigroup as a $(2,1,1)$-type subalgebra is known to be undecidable. In this article, we investigate the problem for certain classes of ample semigroups. We also give examples of semigroups that are left (respectively, right) but not right (respectively, left) ample.
Key words and phrases:
Ample semigroup, Inverse semigroup, Brandt Semigroup, Embedding, Dominion
2010 Mathematics Subject Classification:
20M18, 20M10, 20M15
This research is being supported by the Estonian Research Council grant PRG1204.
1. Motivation and preliminaries
Let $X$ be a (possibly empty) set. By a partial bijection of $X$ we mean a bijection $\beta:A\longrightarrow B$ such that $A$ and $B$ are subsets of $X$, called, respectively, the domain and codomain of $\beta$. We shall denote $A$ and $B$ by $Dom\beta$ and $Im\beta$, respectively. The set of all partial bijections of $X$ is denoted by $\mathcal{I}_{X}$. The empty map (which has empty domain and codomain) may be viewed as a special element of $\mathcal{I}_{X}$. Throughout this article, we shall be following the convention of writing maps to the right of their arguments. Also, parentheses around arguments will be dropped if there is no chance of ambiguity. Given $\beta,\gamma\in\mathcal{I}_{X}$, we define
$\beta\gamma=\beta|_{(Im\beta\cap Dom\gamma)\beta^{-1}}\circ\gamma,$ (1.1)
where $\circ$ denotes the ordinary composition of (full) maps. The binary operation defined by (1.1) turns $\mathcal{I}_{X}$ into a semigroup, called the symmetric inverse semigroup over $X$. Because the empty map acts as the zero element of $\mathcal{I}_{X}$, we shall denote it by 0. Note that for every $\beta\in\mathcal{I}_{X}$ there exists a unique partial bijection $\beta^{-1}\in\mathcal{I}_{X}$, such that $\beta\beta^{-1}$ and $\beta^{-1}\beta$ are identities on $Dom\beta$ and $Im\beta$, respectively. Particularly, $\beta\beta^{-1}\beta=\beta$ and $\beta^{-1}\beta\beta^{-1}=\beta^{-1}$.
In general, an element $a$ of a semigroup $S$ is said to be invertible if there exists a unique element $a^{-1}\in S$, called the inverse of $a$, such that $aa^{-1}a=a$ and $a^{-1}aa^{-1}=a^{-1}$. We call $S$ an inverse semigroup if all of its elements are invertible. The semigroup $\mathcal{I}_{X}$, discussed above, is a classical example of an inverse semigroup. From now on, the letter $T$, possibly with subscripts, will be reserved to only denote inverse semigroups.
A morphism $h:T_{1}\longrightarrow T_{2}$ of inverse semigroups is just a semigroup homomorphism from $T_{1}$ to $T_{2}$. Monomorphisms and isomorphisms of (inverse) semigroups are, respectively, the injective and bijective homomorphisms. The Wagner-Preston representation (see Theorems 1.1 and 1.2) asserts that every inverse semigroup can be embedded in a symmetric inverse semigroup.
Theorem 1.1.
Given an inverse semigroup $T$ the map
$\rho_{x}:Txx^{-1}\longrightarrow Tx^{-1}x,$
defined by
$(a)\rho_{x}=ax,\,\,\forall\,a\in Txx^{-1},$
is a bijection for all $x\in T$. Furthermore, the map
$\rho_{{}_{T}}:T\longrightarrow\mathcal{I}_{T},\text{ given by }\,x\longmapsto\rho_{x},\,\,\forall\,x\in T,$
is a monomorphism.
Proof.
See, for instance, [5] Theorem 5.1.7. ∎
Theorem 1.1 may also be presented in the following dual form.
Theorem 1.2.
Let $T$ be an inverse semigroup. Then, the map
${\lambda_{x}}:x^{-1}xT\longrightarrow xx^{-1}T,$
given by
$(a){\lambda_{x}}=xa,\,\,\forall\,a\in x^{-1}xT,$
is a bijection for all $x\in T$. Moreover, the map
$\lambda_{T}:T\longrightarrow\mathcal{I}_{T},\text{ defined as }\,x\longmapsto{\lambda_{x}},\,\,\forall\,x\in T,$
is a monomorphism.
Proof.
See, for instance, [8] Theorem 1.5.1. ∎
Clearly, for any inverse semigroup $T$ the map $\vartheta:(T)\rho_{{}_{T}}\longrightarrow(T)\lambda_{T}$, defined by $\rho_{x}\mapsto{\lambda_{x}}$, for all $x\in T$, is an isomorphism, where $\rho_{{}_{T}}$ and $\lambda_{T}$ are the representations given by Theorems 1.1 and 1.2, respectively. Also,
[…]
As usual, the set of idempotents of a semigroup $S$ will be denoted by $E(S)$. For an inverse semigroup $T$, we have
$E(T)=\{aa^{-1}:a\in T\};$
moreover, the latter is a subsemilattice of the former. Identifying $T$ with its isomorphic copy in $\mathcal{I}_{T}$ we see that the idempotents $xx^{-1},x^{-1}x\in E(T)$ are, respectively, the identities on the ‘domain’ and ‘codomain’ of $x\in T$. This fact may be used to verify that
$\forall\,x,y\in T,\,x\leq y\text{ iff }x=xx^{-1}y$ (1.3)
defines a partial order on TT—interpret (1.3) as $x\leq y$ if and only if the domain of $x$ is a subset of the domain of $y$ and the latter agrees with the former on the restricted domain. It is an easy exercise to show that Condition (1.3) is equivalent to the following one:
$\forall\,x,y\in T,\,x\leq y\text{ iff }\,\exists\,e\in E(T),\text{ such that }x=ey.$ (1.4)
We call $\leq$ the natural partial order on $T$. In fact, for any semigroup $S$ the set $E(S)$ comes equipped with the partial order:
$e\preccurlyeq f\,\text{ iff }\,ef=fe=e,\;\forall\,e,f\in E(S).$ (1.5)
[…]
A semigroup $S$ is called left ample if it can be embedded in an inverse semigroup $T$ such that $(x\phi)(x\phi)^{-1}\in S\phi$ for all $x\in S$, where $\phi$ is the embedding of $S$ into $T$. Identifying $S$ with its isomorphic copy in $T$, we shall henceforth write $xx^{-1}=[(x\phi^{\prime})(x\phi^{\prime})^{-1}]\phi^{\prime-1}$, where $\phi^{\prime}:S\longrightarrow(S)\phi$ is the isomorphism defined by $x\mapsto(x)\phi$ for all $x\in S$. So, we can say that a left ample semigroup comes equipped with a unary operation $x\mapsto x^{+}$, defined by $x^{+}=xx^{-1}$. Similarly, a semigroup $S$ is called right ample if there exists an inverse semigroup $T$ admitting a monomorphism $\phi:S\longrightarrow T$ such that $(x\phi)^{-1}(x\phi)\in S\phi$. We then define $x^{-1}x=\left[(x\phi^{\prime})^{-1}(x\phi^{\prime})\right]\phi^{\prime-1}\in S$ for all $x\in S$. A right ample semigroup comes endowed with the unary operation $x\mapsto x^{*}$, given by $x^{*}=x^{-1}x$. We shall refer to $T$ as an associated inverse semigroup of the (right, left) ample semigroup $S$. One may easily verify that $S$ is left (respectively, right) ample in $T$ if and only if $(S)\rho_{{}_{T}}$ is left (respectively, right) ample in $\mathcal{I}_{T}$, where $\rho_{{}_{T}}$ is the monomorphisms given by Theorem 1.1. This is also true for $(S)\lambda_{T}$, with $\lambda_{T}$ being the monomorphism given by Theorem 1.2. Note that, a left or a right ample semigroup may be viewed as an algebra of type $(2,1)$. A semigroup $S$ is called ample if it has (possibly different) associated inverse semigroups $T_{1}$ and $T_{2}$ making it, respectively, into a right and a left ample semigroup. An ample semigroup may be considered a $(2,1,1)$-type algebra. If $S$ is a left (respectively, right) ample subsemigroup of an associated inverse semigroup $T$ then we shall say that $S$ is left, (respectively, right) ample in $T$. Similarly, $S$ will be called ample in $T$ if it is both left and right ample in $T$. Obviously, every inverse semigroup is ample (in itself), and is, as such, a $(2,1,1)$-type algebra. Also, a subsemigroup $S$ of an inverse semigroup $T$ is ample in $T$ if $E(T)\subseteq S$, that is, if $S$ is full in $T$. The converse is not true; for example, $\mathbb{N}$ is ample but not full in the multiplicative monoid $\mathbb{Q}$. For more examples of (left, right) ample semigroups the reader is referred to [3] and [6] and the references contained therein.
Problem 1.3.
Let $T_{1}$ and $T_{2}$ be different inverse semigroups containing isomorphic copies, say $S_{1}$ and $S_{2}$, of a non-inverse semigroup $S$. Let $S_{1}$ be left ample in $T_{1}$ and $S_{2}$ be right ample in $T_{2}$. Then, can we find an (associated) inverse semigroup $T$ making (an isomorphic copy of) $S$ into a left as well as right ample semigroup? In the language of universal algebra, this amounts to embedding $S$ in an inverse semigroup as a $(2,1,1)$-type subalgebra.
The following results imply that, in general, Problem 1.3 is undecidable even for finite ample semigroups.
Theorem 1.4 ([4], Theorem 3.4).
Let $S$ be an ample semigroup. Then $S$ is a $(2,1,1)$-type subalgebra of an inverse semigroup if and only if $S$ is a full subsemigroup of an inverse semigroup.
Corollary 1.5 ([4], Corollary 4.3).
It is undecidable whether a finite ample semigroup embeds as a full subsemigroup of a finite inverse semigroup, or of an inverse semigroup.
In Section 2, we shall consider Problem 1.3 for certain classes of (finite) semigroups. Also, to the knowledge of the authors, there exists no example of a left (respectively, right) ample semigroup that is not right (respectively, left) ample. A class of such semigroups will be provided in Section 3. We shall also consider Problem 1.3 for (left, right) rich ample semigroups that were introduced in [6]. For the convenience of the reader, their definitions have been reproduced below.
For a subsemigroup $S$ of an inverse semigroup $T$ let us define $S^{\prime}=\{s^{-1}\in T:s\in S\}$. We say that $S$ is rich left ample (respectively, rich right ample) in $T$ if for all $x,y\in S$ one has $xy^{-1}\in S\cup S^{\prime}$ (respectively, $x^{-1}y\in S\cup S^{\prime}$). Alternatively (see [6], Proposition 4.8), $S$ is rich left ample in $T$ if
  1. (1L)
    $S$ is left ample in $T$, and
  2. (2L)
    for all $x,y\in S$, $xy^{-1}=ayy^{-1}$, where $a\in S\cup S^{\prime}$.
Similarly, the conjunction of,
  1. (1R)
    $S$ is right ample in $T$, and
  2. (2R)
    for all $x,y\in S$, $x^{-1}y=x^{-1}xb$, where $b\in S\cup S^{\prime}$,
is equivalent to $S$ being rich right ample in $T$ ([6], Proposition 4.9). Obviosly, in Conditions (2L) and (2R), one may replace $yy^{-1}$ and $x^{-1}x$ by $y^{+}$ and $x^{*}$, respectively. We call $S$ rich left (right) ample if it can be embedded, as such, in an (associated) inverse semigroup. We say that $S$ is (two-sided) rich ample if it is rich left as well as rich right ample—again, the associated inverse semigroups need not be the same. If the element $a$ in Condition (2L) is determined uniquely then $S$ is said to be ultra rich left ample. Ultra rich right and ultra rich ample semigroups are defined analogously.
Remark 1.6.
Let a semigroup $S$ be left (respectively, right) ample in an inverse semigroup $T_{1}$, and $T_{2}$ be an inverse semigroup admitting a morphism $f:T_{1}\longrightarrow T_{2}$. Then one can easily verifiy that $(S)f$ is left (respectively, right) ample in $T_{2}$. The remark also holds for rich (left, right) ample semigroups. However, ultra rich (left, right) ample subsemigroups are only preserved by monomorphisms.
From any ample semigroup $S$ one may obtain an amalgam $(S;T_{1},T_{2};\phi_{1},\phi_{2})$, see [6] for details, where $T_{i},1\leq i\leq 2,$ are (associated) inverse semigroups admitting monomorphisms $\phi_{i}:S\longrightarrow T_{i}$ such that $S_{1}=(S)\phi_{1}$ is left ample in $T_{1}$ and $S_{2}=(S)\phi_{2}$ is right ample in $T_{2}$. Let $V_{1}$ and $V_{2}$ be the inverse hulls of $S_{1}$ and $S_{2}$ in $T_{1}$ and $T_{2}$, respectively (see [6]). Also, assume that the isomorphism $\psi=\phi_{1}^{-1}\circ\,\phi_{2}:(S)\phi_{1}\longrightarrow(S)\phi_{2}$ extends to a homomorphism, say $\psi^{\prime}$, from $V_{1}$ to $V_{2}$. Now, because every element of $V_{1}$ may be written as $x_{1}x_{2}\dots x_{n}$, where for each $i\in\{1,2,\dots,n\}$ either $x_{i}$ or $x_{i}^{-1}$ belongs to $S_{1}$, $\psi^{\prime}$ must be defined by
$(x_{1}x_{2}\dots x_{n})\psi^{\prime}=(x_{1})\overline{\psi}(x_{2})\overline{\psi}\dots(x_{n})\overline{\psi},$ (1.6)
where
[…]
Because $(V_{1})\psi^{\prime}$ is an inverse subsemigroup of $V_{2}$ containing $S_{2}$, we may further assert that $\psi^{\prime}$ is necessarily surjective. This implies by Remark 1.6 that $S_{2}=(S_{1})\psi^{\prime}$ is right as well as left ample in $V_{2}$, and hence in $T_{2}$. Furthermore, there exists a (retract) morphism $\psi^{\prime\prime}:V_{2}\longrightarrow V_{1}$ such that $\psi^{\prime\prime}\circ\psi^{\prime}$ is identity on $V_{2}$, see, for instance, [9] §5.6. Because, by a similar argument, $\psi^{\prime\prime}$ is also surjective, it follows that $\psi^{\prime\prime}=(\psi^{\prime})^{-1}$, implying that $\psi^{\prime}$ is an isomorphism. This implies that $S_{1}$ is also left as well as right ample in $V_{1}$, equivalently, in $T_{1}$. Condition (1) of the following proposition recapitulates the above discussion.
Theorem 1.7.
Let $T_{i},1\leq i\leq 2$, be inverse semigroups admitting monomorphisms $\phi_{i}:S\longrightarrow T_{i}$ from a semigroup $S$. Suppose that $(S)\phi_{1}$ is left ample in $T_{1}$ and $(S)\phi_{2}$ is right ample in $T_{2}$. Then $(S)\phi_{i}$ is ample in $T_{i}$ for each $i\in\{1,2\}$ if one of the following (equivalent) conditions is satisfied. Particularly, $(S)\phi_{i}$ is a $(2,1,1)$-type subalgebra of $T_{i}$, for each $i\in\{1,2\}$.
  1. (1)
    The isomorphism $\phi_{1}^{-1}\circ\phi_{2}:(S)\phi_{1}\longrightarrow(S)\phi_{2}$ can be extended to a homomorphism (equivalently, isomorphism) from the inverse hull of $(S)\phi_{1}$ in $T_{1}$ to the inverse hull of $(S)\phi_{2}$ in $T_{2}$.
  2. (2)
    If $x_{1}x_{2}\dots x_{n}=y_{1}y_{2}\dots y_{m}$ in $V_{1}$ then $(x_{1})\overline{\psi}(x_{2})\overline{\psi}\dots(x_{n})\overline{\psi}=(y_{1})\overline{\psi}(y_{2})\overline{\psi}\dots(y_{m})\overline{\psi}$ in $V_{2}$, where $V_{1}$, $V_{2}$ and $\overline{\psi}$ are as defined above.
  3. (3)
    The amalgam $(S;T_{1},T_{2};\phi_{1},\phi_{2})$ is weakly embeddable in an inverse semigroup (see [6] for definition of weak embedding).
Proof.
We only need to show that the three conditions are equivalent.
$(1)\Longleftrightarrow(3)$ follows from from [6] Theorem 4.2.
$(1)\Longrightarrow(2)$ is obvious.
As for the implication $(2)\Longrightarrow(1)$, observe that Condition (2) precisely says that the map $\psi^{\prime}:V_{1}\longrightarrow V_{2}$ given by (1.6) is well-defined. But then, it also follows from its very definition that $\psi^{\prime}$ is a homomorphism that extends $\phi_{1}^{-1}\circ\phi_{2}$. (That the extension is indeed an isomorphism, has already been observed in the discussion preceding the theorem.) ∎
2. Answering problem 1.3 for certain classes of ample semigroups
Let $S$ be a subsemigroup of an inverse semigroup $T$. Following [1], a subset $Y$ of $T$ will be called right $S$-invariant if
$Im(\rho_{s}|_{Dom\rho_{s}\cap Y})\subseteq Y,\hskip 4.0pt\text{for all }s\in S,$
where $\rho_{s}:Tss^{-1}\longrightarrow Ts^{-1}s$ is the partial bijection defined in Theorem 1.1. Denoting $\rho_{s}|_{Dom\rho_{s}\cap Y}$, equivalently $\rho_{s}|_{Tss^{-1}\cap Y}$, by $\sigma_{s}$, we introduce a map:
$\hat{\rho}_{{}_{Y}}:S\longrightarrow\mathcal{I}_{Y}\text{ by }(s)\hat{\rho}_{{}_{Y}}=\sigma_{s}.$
Clearly, $\hat{\rho}_{{}_{Y}}$ is well-defined. Also, it is known to be a homomorphism when $S$ is finite; see [1]. The following lemma shows that $\hat{\rho}_{{}_{Y}}$ is a homomorphism even if $S$ is not finite.
Lemma 2.1.
With the notations defined above, $\hat{\rho}_{{}_{Y}}$ is a homomorphism.
Proof.
We need to show that $\sigma_{s}\sigma_{t}=\sigma_{st}$, for all $s,t\in S$. By Theorem 1.1 we have $\rho_{st}=\rho_{s}\rho_{t}$, also requiring that
$Dom\rho_{st}=Dom(\rho_{s}\rho_{t})=(Ts^{-1}s\cap Ttt^{-1})\rho^{-1}_{s}.$
Consequently,
$\sigma_{st}=\rho_{st}|_{Dom\rho_{st}\cap Y}=(\rho_{s}\rho_{t})|_{(Ts^{-1}s\cap Ttt^{-1})\rho^{-1}_{s}\cap Y}.$
On the other hand, one has
$\sigma_{s}\sigma_{t}=\left(\rho_{s}|_{Tss^{-1}\cap Y}\right)\left(\rho_{t}|_{Ttt^{-1}\cap Y}\right).$
whence,
$(x)\sigma_{st}=(x)(\rho_{s}\rho_{t})=(x)(\sigma_{s}\sigma_{t}),\,\forall\,x\in Dom(\sigma_{s}\sigma_{t})\cap Dom(\sigma_{st}).$
So, the proof will be accomplished if we show that $Dom(\sigma_{s}\sigma_{t})=Dom(\sigma_{st})$, that is,
$(Tss^{-1}\cap Y)\rho_{s}\cap(Ttt^{-1}\cap Y)]\rho^{-1}_{s}=(Ts^{-1}s\cap Ttt^{-1})\rho^{-1}_{s}\cap Y.$
To this end, observe that
[…]
where the ‘if’ part of the last implication follows by observing that
$x\in(Ts^{-1}s\cap Ttt^{-1})\rho^{-1}_{s}\implies x\rho_{s}\in Ts^{-1}s\implies x\in Tss^{-1}\implies x\in Dom\rho_{s}.$
This implies that $Dom(\sigma_{s}\sigma_{t})=Dom(\sigma_{st})$, and the proof is complete. ∎
In particular, the subsemigroup $S$ is a right $S$-invariant subset of $T$, giving a homomorphism $\hat{\rho}_{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$, as defined above.
Proposition 2.2.
If $S$ is left ample in $T$, then $\hat{\rho}_{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$ is a monomorphism. Moreover, $(S)\hat{\rho}_{{}_{S}}$ is left ample in $\mathcal{I}_{S}$.
Proof.
To prove the first part of the proposition, we only need to show that $\hat{\rho}_{{}_{S}}$ is injective. Assume, for this purpose, that
$\sigma_{s}=(s)\hat{\rho}_{{}_{S}}=(t)\hat{\rho}_{{}_{S}}=\sigma_{t},\;\text{i.e. }\,\rho_{s}|_{Tss^{-1}\cap S}=\rho_{t}|_{Ttt^{-1}\cap S},$
for some $s,t\in S$. Then, $ss^{-1}\in Tss^{-1}\cap S$ and $tt^{-1}\in Ttt^{-1}\cap S$, because $S$ is left ample, and we may calculate from $Tss^{-1}\cap S=Ttt^{-1}\cap S$ (viz. the equality of domains of $\sigma_{s}$ and $\sigma_{t}$) that
$(ss^{-1})\rho_{s}=(ss^{-1})\rho_{t}\;\text{ and }\;(tt^{-1})\rho_{s}=(tt^{-1})\rho_{t}.$
This gives
$s=ss^{-1}s=ss^{-1}t\;\text{ and }\;t=tt^{-1}t=tt^{-1}s.$
Consequently, we have $s\leq t\leq s$ in $T$, yielding $s=t$. Hence, $\hat{\rho}_{{}_{S}}$ is an injection.
As for the second part of the proposition, let us first notice that
$Sxx^{-1}\subseteq Txx^{-1}\cap S,\ \forall\,x\in S.$
Then observe that, for all $y\in T,\,x\in S$, we have
$y(xx^{-1})\in Txx^{-1}\cap S\implies y(xx^{-1})=y(xx^{-1})(xx^{-1})\in Sxx^{-1}.$
Thus, for every $x\in S$, one has
$Sxx^{-1}=Txx^{-1}\cap S.$ (2.1)
This implies that the domain and codomain of $\sigma_{s}$ are indeed
$Sss^{-1}\text{ and }Sss^{-1}s=Ss\,\,(\subseteq Ss^{-1}s\cap S),$ (2.2)
respectively. Let us next note that
$(\rho_{{}_{T}})^{-1}\circ\hat{\rho}_{{}_{S}}:(S)\rho_{{}_{T}}\longrightarrow(S)\hat{\rho}_{{}_{S}}$
defined by
$\rho_{s}\longmapsto\rho_{s}|_{Dom\rho_{s}\cap S}=\sigma_{s},\,s\in S,$
is an isomorphism. Now, because $\sigma_{x}\sigma_{x}^{-1}$ is the identity on $Sxx^{-1}=Txx^{-1}\cap S$ and $\rho_{xx^{-1}}$ is the identity on $Txx^{-1}$ ($Dom\rho_{x}$), where $x^{-1}$ is the inverse of $x$ in $T$, we have
$\sigma_{x}\sigma_{x}^{-1}=\rho_{xx^{-1}}|_{Dom\rho_{xx^{-1}}\cap S}.$
Finally, since $xx^{-1}\in S$, it follows from the above equality that
$\sigma_{x}\sigma_{x}^{-1}=(\rho_{xx^{-1}})\left((\rho_{{}_{T}})^{-1}\circ\hat{\rho}_{{}_{S}}\right)=\left((\rho_{xx^{-1}})(\rho_{{}_{T}})^{-1}\right)\hat{\rho}_{{}_{S}}\in(S)\hat{\rho}_{{}_{S}}.$
Hence $\sigma_{x}\sigma_{x}^{-1}\in(S)\hat{\rho}_{{}_{S}}$, implying that $(S)\hat{\rho}_{{}_{S}}$ is left ample in $\mathcal{I}_{S}$. ∎
Corollary 2.3.
Let $S$ be left ample in $T$. Also, let $\hat{\rho}_{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$ and $\sigma_{s}:Sss^{-1}\longrightarrow Ss$ be as considered in Proposition 2.2. Then, considering the isomorphism
$\theta=(\rho_{{}_{T}}|_{S})^{{}^{-1}}\circ\hat{\rho}_{{}_{S}}:(S)\rho_{{}_{T}}\longrightarrow\mathcal{I}_{S},\ \text{ defined by }\ \rho_{s}\longmapsto\sigma_{s},\forall\,s\in S,$
we have, for all $s\in S$,
$(\rho_{s}\rho_{s}^{-1})\theta=\sigma_{s}\sigma_{s}^{-1}=\sigma_{ss^{-1}},$
where $ss^{-1}=(\rho_{s}\rho_{s}^{-1})\rho_{{}_{T}}^{-1}$.
Proof.
Using (2.1), let us first calculate
$\sigma_{ss^{-1}}=\rho_{ss^{-1}}|_{Sss^{-1}}=(\rho_{s}\rho_{s^{-1}})|_{Sss^{-1}}=(\rho_{s}\rho_{s}^{-1})|_{Sss^{-1}}=(\rho_{s}|_{Sss^{-1}})(\rho_{s}|_{Sss^{-1}})^{-1}=\sigma_{s}\sigma_{s}^{-1}.$
Then, we have
$(\rho_{s}\rho_{s}^{-1})\theta=(\rho_{s}\rho_{s^{-1}})\theta=(\rho_{ss^{-1}})\theta=\left((\rho_{ss^{-1}})(\rho_{{}_{T}}|_{S})^{{}^{-1}}\right)\hat{\rho}_{{}_{S}}=(ss^{-1})\hat{\rho}_{{}_{S}}=\sigma_{ss^{-1}}=\sigma_{s}\sigma_{s}^{-1},$
completing the proof. ∎
For any left ample semigroup $S$, it follows from the above corollary that for all $x\in S$,
$\left(\sigma_{x}\sigma_{x}^{-1}\right)\hat{\rho}_{{}_{S}}^{-1}=\left(\rho_{x}\rho_{x}^{-1}\right){\rho}_{{}_{T}}^{-1}=xx^{-1}\in S.$
The following theorem is essentially a restatement of Proposition 2.2.
Theorem 2.4.
Let $U$ be a left ample semigroup (with some associated inverse semigroup $T$). Then $\hat{\rho}_{{}_{U}}:U\longrightarrow\mathcal{I}_{U}$, given by $x\mapsto\sigma^{\prime}_{x}$, where $\sigma^{\prime}_{x}:Uxx^{-1}\longrightarrow Ux$ is defined by $(z)\sigma^{\prime}_{x}=zx$ for all $z\in Uxx^{-1}$, is a monomorphism. Moreover, $(U)\hat{\rho}_{{}_{U}}$ is left ample in $\mathcal{I}_{U}$.
Proof.
The morphisms maps used in the proof are shown in Figure 1 (the morphism $\theta$ was defined in Corollary 2.3). Let $T$ be an (associated) inverse semigroup admitting a monomorphism
$\phi:U\longrightarrow T,$
such that $(U)\phi=S$ is left ample in $T$. Let $\phi^{\prime}:U\longrightarrow S$ be the isomorphism defined by $x\mapsto(x)\phi$, and $\hat{\rho}_{{}_{S}}:S\longrightarrow\mathcal{I}_{S}$, defined by
$(x)\phi^{\prime}\longmapsto\sigma_{(x)\phi^{\prime}},\,\forall\,x\in U,$
be the embedding given by Proposition 2.2. Also, let $\hat{\phi}=\phi^{\prime}\circ\hat{\rho}_{{}_{S}}:U\longrightarrow\mathcal{I}_{S}$. Then, defining $s=(x)\phi^{\prime}$, and using Corollary 2.3, we obtain
[…]
where the last equality is due to the identification: $(x)\phi^{\prime}((x)\phi^{\prime})^{-1}=(xx^{-1})\phi^{\prime}$. Thus, we may write
$\{(x)\hat{\phi}[(x)\hat{\phi}]^{-1}\}(\hat{\phi})^{-1}=xx^{-1}=\{[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}\}(\phi^{\prime})^{-1}\in U.\vskip 2.0pt$ (2.4)
Figure 1.
Now, for an arbitrary partial bijection $\beta:A\longrightarrow B$ in $\mathcal{I}_{S}$, consider the map
$\gamma:(A)(\phi^{\prime})^{-1}\longrightarrow(B)(\phi^{\prime})^{-1}$
defined by
$(a)(\phi^{\prime})^{-1}\longmapsto((a)\beta)(\phi^{\prime})^{-1},\,\forall\,a\in A.$ (2.5)
Clearly, $\gamma$ is an element of $\mathcal{I}_{U}$, and it is an easy exercise to prove that $\chi:\mathcal{I}_{S}\longrightarrow\mathcal{I}_{U}$ given by $\beta\mapsto\gamma$ is an isomorphism. Moreover, by Remark 1.6, $((U)\hat{\phi})\chi$ is left ample in $\mathcal{I}_{U}$, proving the second part of the proposition.
To complete the proof, let us consider the monomorphism
$\pi=\hat{\phi}\circ{\chi}\colon U\longrightarrow\mathcal{I}_{U}.$
The aim is to show that for all $x\in U$, $(x)\pi=\sigma^{\prime}_{x}$ where $\sigma^{\prime}_{x}$ is as given in the statement of the theorem. Recall that the domain and codomain of $\sigma_{(x)\phi}$ ($\sigma_{s}$) are, respectively, $(S)[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}$ and $(S)[(x)\phi^{\prime}]$. Hence, by the definitions of $\pi$ and $\chi$, we have
$(x)\pi=\left(\sigma_{(x)\phi}\right)\chi:\{S[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}\}(\phi^{\prime})^{-1}\longrightarrow\{S[(x)\phi^{\prime}]\}(\phi^{\prime})^{-1},\vskip 2.0pt$
defined according to (2.5). However, it is immediate that
$\{S[(x)\phi^{\prime}]\}(\phi^{\prime})^{-1}=Ux,$
and by (2.4), we have $\{[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}\}(\phi^{\prime})^{-1}=xx^{-1}$, whence
$\{S[(x)\phi^{\prime}][(x)\phi^{\prime}]^{-1}\}(\phi^{\prime})^{-1}=Uxx^{-1}.$
Also, taking $z=uxx^{-1},u\in U$, we may calculate,
$\begin{split}(z)[(x)\pi]&=((uxx^{-1})[(x)\hat{\phi}\circ\chi]\\ &=\left(uxx^{-1}\right)\left(\left(\sigma_{(x)\phi^{\prime}}\right)\chi\right)\\ &=((uxx^{-1})\phi^{\prime}(x)\phi^{\prime})\phi^{\prime-1}\\ &=((uxx^{-1}x)\phi^{\prime})\phi^{\prime-1}\\ &=uxx^{-1}x\\ &=zx.\end{split}$
So, we may conclude that
$(x)\pi:Uxx^{-1}\longrightarrow Ux,$
is defined by $(z)[(x)\pi]=zx$ for all $z=uxx^{-1}\in Uxx^{-1}$, where $u\in U$. ∎
Theorem 2.5.
Let $S$ be left as well as right ample in $T$ with $Ts^{-1}s\cap S=Ss$ for all $s\in S$. Then $(S)\hat{\rho}_{{}_{S}}$ is left as well right ample in $\mathcal{I}_{S}$.
Proof.
Using the right ampleness of $S$ we have
$(s^{-1}s)\rho_{{}_{T}}=\rho_{s^{-1}s}=\rho_{{}_{s}}^{-1}\rho_{{}_{s}}.$
Because $\rho_{{}_{s}}^{-1}\rho_{{}_{s}}$ is identity on $Ts^{-1}s$, it follows from the definition of $\theta$ that
$(s^{-1}s)\hat{\rho}_{{}_{S}}=\left(\rho_{s^{-1}s}\right)\theta=\sigma_{s^{-1}s}$
is identity on $Ts^{-1}s\cap S$.
On the other hand $\sigma_{s^{-1}}\sigma_{s}$ is identity on $Ss$, which by the assumption is equal to $Ts^{-1}s\cap S$. Thus, we have
$(s^{-1}s)\hat{\rho}_{{}_{S}}=\left(\rho_{s^{-1}s}\right)\theta=\sigma_{s^{-1}s}=\sigma_{s^{-1}}\sigma_{s},$
implying that $(S)\hat{\rho}_{S}$ is right ample in $\mathcal{I}_{S}$. That $(S)\hat{\rho}_{S}$ is left ample in $\mathcal{I}_{S}$ follows from Theorem 2.4. ∎
Problem 2.6.
What is the situation of Theorem 2.5 with rich ample semigroups?
Similarly, using the dual notion of left $S$-invariance, one can prove that any right ample subsemigroup $S$ of an inverse semigroup $T$ can be embedded in the symmetric inverse semigroup $\mathcal{I}_{S}$ via the monomorphism,
$\hat{\lambda}_{S}:S\longrightarrow\mathcal{I}_{S},\hskip 5.0pt\text{given by }s\longmapsto{{}_{s}\sigma}={\lambda_{s}}|_{s^{-1}sT\,\cap\,S},\,\forall\,s\in S,$
where $\lambda_{s}\in\mathcal{I}_{T}$ is the partial bijection defined in Theorem 1.2. Because, in this case, $s^{-1}sT\cap S=s^{-1}sS$, we have the following analogue of Theorem 2.4.
Theorem 2.7.
Any right ample semigroup $U$ may be embedded in the symmetric inverse semigroup $\mathcal{I}_{U}$ via,
$\hat{\lambda}_{U}:U\longrightarrow\mathcal{I}_{U},\text{ given by }u\longmapsto{{}_{u}\sigma^{\prime}},$
where the partial bijection,
${{}_{u}\sigma^{\prime}}:u^{-1}uU\longrightarrow uU$
is given by $(x){{}_{u}\sigma^{\prime}}=ux$, for all $x\in u^{-1}uU$. Also, $(U){\hat{\lambda}_{U}}$ is right ample in $\mathcal{I}_{U}$.
Combining Theorems 2.4 and 2.7, we see that any ample semigroup $S$ has two isomorphic copies, viz. $(S)\hat{\rho}_{{}_{S}}$ and $(S)\hat{\lambda}_{S}$, in $\mathcal{I}_{S}$ that are, respectively, left and right ample in $\mathcal{I}_{S}$. This also gives the amalgam $\mathcal{A}=(S;\mathcal{I}_{S},\mathcal{I}_{S};\hat{\rho}_{{}_{S}},\hat{\lambda}_{S})$ (see, for instance, [6] for definition). Note that $\mathcal{A}$ is not known to be a special amalgam, because the isomorphism
$\left(\hat{\rho}_{{}_{S}}\right)^{-1}\circ\left(\hat{\lambda}_{S}\right):(S)\hat{\rho}_{{}_{S}}\longrightarrow{(S)\hat{\lambda}_{S}}$
may not extend to an automorphism of $\mathcal{I}_{S}$.
Theorem 2.8.
An ample semigroup $S$ is a $(2,1,1)$-type subalgebra of $\mathcal{I}_{S}$ if one of the following (equivalent) conditions is satisfied.
  1. (1)
    The isomorphism $\psi=(\hat{\rho}_{{}_{S}})^{-1}\circ\,\hat{\lambda}_{S}:(S)\hat{\rho}_{{}_{S}}\longrightarrow(S)\hat{\lambda}_{S}$ can be extended to a homomorphism (equivalently, isomorphism) from $V_{1}$ to $V_{2}$, that are the inverse hulls of $(S)\hat{\rho}_{{}_{S}}$ and $(S)\hat{\lambda}_{S}$, respectively.
  2. (2)
    If $x_{1}x_{2}\dots x_{n}=y_{1}y_{2}\dots y_{m}$ in $V_{1}$ then $(x_{1})\overline{\psi}(x_{2})\overline{\psi}\dots(x_{n})\overline{\psi}=(y_{1})\overline{\psi}(y_{2})\overline{\psi}\dots(y_{m})\overline{\psi}$ in $V_{2}$, where $V_{1}$ and $V_{2}$ are the inverse hulls mentoined in (1), and $\overline{\psi}$ is given by (1.6).
  3. (3)
    The amalgam $\mathcal{A}=(S;\mathcal{I}_{S},\mathcal{I}_{S};\hat{\rho}_{{}_{S}},\hat{\lambda}_{S})$ is weakly embeddable.
Proof.
Similar to the proof of Theorem 1.7. ∎
Remark 2.9.
The converse of Theorem 2.8 is not true. To see this, consider an arbitrary finite ample semigroup $S$ (see [7] for examples). Now, if the converse of Theorem 2.8 were true, then each of its Conditions (1) and (2) would be necessary and sufficient for $S$ to be a $(2,1,1)$-subalgebra of an inverse semigroup. This contradicts the undecidability of Problem 1.3 for finite semigroups (cf. Theorem 1.4 and Corollary 1.5) because the truth of Conditions (1) and (2) could be determined in finitely many steps.
Example 2.10.
An example is needed where a finite ample semigroup $S$ is a $(2,1,1)$-subalgebra of an inverse semigroup $T$ but is not such in $\mathcal{I}_{S}$.
Theorem 2.11.
Let $S$ be an ample semigroup such that for all $e\in E(S)$ there exists a bijection between the principal ideals $eS$ and $Se$. Then $S$ is ample in (and hence a $(2,1,1)$-subalgebra of) $\mathcal{I}_{S}$.
Proof.
If $S$ is an inverse semigroup then there is nothing to prove. So, assume that $S$ is non-inverse. Let $S$ be left ample in $T_{1}$ and right ample in $T_{2}$. Let $x$ be an arbitrary element of $S$ and $\phi_{i},1\leq i\leq 2,$ denote the embeddings of $S$ in $T_{i}$. Then
$(x)\phi_{1}((x)\phi_{1})^{-1}\in(S)\phi_{1},\text{ and }\,((x)\phi_{2})^{-1}(x)\phi_{2}\in(S)\phi_{2}.$
Let
$y=((x)\phi_{1}((x)\phi_{1})^{-1})\phi_{1}^{-1},\text{ and }\,z=(((x)\phi_{2})^{-1}(x)\phi_{2})\phi_{2}^{-1}.$
Then, clearly $y,z\in E(S)$.
Next, by Theorems 2.4 and 2.7, $\hat{S}_{1}=(S)\hat{\rho}_{{}_{S}}$ and $\hat{S}_{2}=(S)\hat{\lambda}_{S}$ are, respectively, left and right ample in $\mathcal{I}_{S}$, with
$(x)\hat{\rho}_{{}_{S}}=\sigma_{x}:Sy\longrightarrow Sx$
and
$(x)\hat{\lambda}_{S}={{}_{x}{\sigma}}:zS\longrightarrow xS.$
Furthermore, from Figure 1, one may assert that
$e_{1}=(y)\hat{\rho}_{{}_{S}}=\sigma_{y}=\sigma_{x}\sigma_{x}^{-1}:Sy\longrightarrow Sy,\hskip 10.0pt\text{ for }\,yy^{-1}=y$
and
$\begin{split}e^{\prime}_{2}=(z)\hat{\lambda}_{S}={{}_{z}\sigma}&={{}_{z}\sigma}:zS\longrightarrow zS,\hskip 40.0pt\text{ since }\,z^{-1}z=z\\ \vskip 4.0pt&={{}_{x}{\sigma}}^{-1}{{}_{x}{\sigma}}:xS\longrightarrow xS\end{split}$
Let us also observe that
$(e^{\prime}_{2})\psi^{-1}=(z)\hat{\rho}_{{}_{S}}=\sigma_{z}:Sz\longrightarrow Sz$
is an idempotent in $(S)\hat{\rho}_{{}_{S}}$, and
$(e_{1})\psi=(y){{}_{S}\hat{\theta}}={{}_{y}\sigma}:yS\longrightarrow yS$
belongs to $E((S)\hat{\lambda}_{S})$. Let
$e^{\prime}_{1}=\sigma^{-1}_{x}\sigma_{x}:Sx\longrightarrow Sx,\;\;\text{(which may not belong to }(S)\hat{\rho}_{{}_{S}}),$
and
$e_{2}={{}_{x}\sigma}{{}_{x}\sigma}^{-1}:zS\longrightarrow zS,\;\;\text{(which may not be a member of }(S)\hat{\lambda}_{S}).$
We may now calculate
$\begin{split}\sigma_{x}\left((e^{\prime}_{2})\psi^{-1}\right)&=\sigma_{x}\left(({{}_{x}\sigma}^{-1}{{}_{x}\sigma})\psi^{-1}\right)\\ &=\left(({{}_{x}\sigma})\psi^{-1}\right)\left(({{}_{x}\sigma}^{-1}{{}_{x}\sigma})\psi^{-1}\right)\\ &=({{}_{x}\sigma}{{}_{x}\sigma}^{-1}{{}_{x}\sigma})(\psi)^{-1}\\ &=({{}_{x}\sigma})(\psi)^{-1}\\ &=\sigma_{x}.\end{split}$
Because $\sigma_{x}^{-1}\sigma_{x}$ is the minimal idempotent such that $\sigma_{x}(\sigma^{-1}_{x}\sigma_{x})=\sigma_{x}$, we must have
$\sigma^{-1}_{x}\sigma_{x}\leq(e^{\prime}_{2})\psi^{-1}=({{}_{x}\sigma}^{-1}{{}_{x}\sigma})\psi^{-1}=({{}_{z}\sigma})\psi^{-1}=\sigma_{z}.$ (2.6)
By a similar token, we also obtain
${{}_{x}\sigma}{{}_{x}\sigma^{-1}}\leq\,{{}_{y}\sigma}.$ (2.7)
The aim is to prove that $Dom(\sigma^{-1}_{x}\sigma_{x})=Dom(\sigma_{z})$ and $Dom({{}_{x}\sigma}{{}_{x}\sigma^{-1}})=Dom({{}_{y}\sigma})$. To this end, we first note, using inequalities (2.6) and (2.7), respectively, that
$Sx=Dom(\sigma^{-1}_{x}\sigma_{x})\subseteq Dom(\sigma_{z})=Sz$
and
$zS=Dom({{}_{x}\sigma}{{}_{x}\sigma^{-1}})\subseteq Dom({{}_{y}\sigma})=yS.$
Hence, we have, by the hypothesis,
$Sx\subseteq Sz=zS\subseteq yS=Sy.$
Now, because $S$ is finite and $\sigma_{x}:Sy\longrightarrow Sx$ is a bijection, we get
$Sx=Sz=zS=yS=Sy.$
Recalling the definition of the natural partial order, it follows from (2.6) and (2.7) that $e^{\prime}_{1}\in(S)\hat{\rho}_{{}_{S}}$ and $e_{2}\in(S)\hat{\lambda}_{S}$. This completes the proof. ∎
The following corollary is straightforward.
Corollary 2.12.
Let $S$ be a finite ample semigroup such that $E(S)$ is contained in the center of $S$. Then $S$ is ample in (equivalently, a $(2,1,1)$-subalgebra of) $\mathcal{I}_{S}$.
Problem 2.13.
What are the inverse semigroups such that the ample semigroups considered in the above theorem and its corollary are their full semigroups (up to isomorphism)?
3. Examples concerning left (right) ample semigroups
A semigroup $S$ without zero is called simple if it has no proper (two-sided) ideals (recall that groups are precisely the semigroups with no proper one-sided ideals). We say that $S$ is completely simple if it is simple and contains an idempotent which is minimal within the set $E(S)$. If a semigroup $S$ has a zero element 0, then clearly $0\leq e$ for all $e\in E(S)$. In this case, an idempotent $f\in E(S)$ is called primitive if $e\leq f$ implies that $e\in\{0,f\}$ for all $e\in E(S)$. A semigroup $S$ with zero is called 0-simple if $S^{2}\neq{0}$ and 0{0} is the only proper ideal of $S$. (The first condition only serves to exclude the two-element null semigroup, which makes the relevant structure theory somewhat clean.) A semigroup $S$ with zero is said to be completely 0-simple if it is 0-simple and contains a primitive idempotent. Completely (simple) 0-simple semigroups form an important class of (regular) semigroups. They are characterized by the celebrated Rees representation theorems (cf. [5], Theorems 3.2.3 and 3.3.1). In this article we shall consider completely 0-simple inverse semigroups that are known as Brandt semigroups. (Completely simple inverse semigroups are in fact groups.) For the convenience of the reader, we are reproducing below the Rees representation theorem for completely 0-simple semigroups, as well as its version for Brandt semigroups. For further background on Rees matrix semigroups the reader may refer to [2], Chapter 3.
Let $G^{0}$ be the semigroup obtained by externally adjoining a zero element 0 to a group $G$ (we call $G^{0}$ a zero group). Let $I$ and $\Lambda$ be some non-empty (indexing) sets and $P=(p_{\lambda i})$ be a $\Lambda\times I$ matrix over $G^{0}$ such that none of its rows or columns consists entirely of zeros (in the literature such a matrix is called regular). We use $P$ to define an associative binary operation on the set $(I\times G\times\Lambda)\cup\{0\}$, as follows:
[…]
$(i,a,\lambda)0=0(i,a,\lambda)=00=0.$
The set $(I\times G\times\Lambda)\cup\{0\}$ together with the above binary operation is called the $I\times\Lambda$ Rees matrix semigroup over the 0-group $G^{0}$ with sandwich matrix $P$. We denote it by $\mathcal{M}^{0}(G^{0},I,\Lambda,P)$.
Theorem 3.1 (Rees representation theorem; [5], Theorem 3.2.3).
Every completely 0-simple semigroup is isomorphic to some Rees matrix semigroup $\mathcal{M}(G^{0},I,\Lambda,P)$.
To characterize Brandt semigroups as a subclass of Rees matrix semigroups, we consider an $I\times I$ sandwich matrix $\Delta$ that has $1$s on the diagonal and zeros elsewhere. We then have the following theorem.
Theorem 3.2 ([5], Theorem 5.1.8).
Every Brandt semigroup is isomorphic to a Rees matrix semigroup $\mathcal{M}^{0}(G^{0},I,I,\Delta)$.
Since the sandwich matrix $\Delta$ will always be obvious, from now on, the Brandt semigroup $B=\mathcal{M}^{0}(G^{0},I,I,\Delta)$ will be denoted by its universe $(I\times G\times I)\cup\{0\}$. Note that the inverse of $(i,g,j)\in B$ is $(j,g^{-1},i)$, where $g^{-1}$ is the inverse of $g$ in $G$. Also, the non-zero idempotents of $B$ are of the form $(i,1,i),\,i\in I$, where $1$ is the identity of $G$.
It is an easy exercise to verify that a semigroup with zero that has no proper non-zero one sided ideal is a zero group (and hence an inverse semigroup). This implies that commutative 0-simple semigroups, equivalently, the commutative Brandt semigroups, are zero (Abelian) groups. It is also a routine to verify that a Brandt semigroup is a zero group if and only if $I$ is a singleton. In this section we shall consider left (right) ample subsemigroups of a Brandt semigroup, such that the latter is not a zero group. So, we shall only be dealing with non-commutative Brandt semigroups. This also means that the set $I$ will contain more than one element.
Definitions 3.3.
By a strict left ample semigroup we mean a semigroup that is left but not right ample. Strict right ample semigroups are defined dually.
Proposition 3.4.
A semigroup $S$ is a strict left ample if and only if the following conditions hold.
  1. (1)
    $(S)\hat{\rho}_{{}_{S}}$ is left ample in $\mathcal{I}_{S}$.
  2. (2)
    For every inverse oversemigroup $T$ of $S$, either $\hat{\lambda}_{{}_{S}}$, given by Theorem 2.7, is not a monomorphism or $(S)\hat{\lambda}_{{}_{S}}$ is not right ample in $\mathcal{I}_{S}$.
Proof.
($\implies$) Let $S$ be strict left ample. Then, by Theorem 2.4, $(S)\hat{\rho}_{{}_{S}}$ is left ample in $\mathcal{I}_{S}$. Also, if $\hat{\lambda}_{S}$ is an isomorohism and $(S)\hat{\lambda}_{S}$ is right ample in $\mathcal{I}_{S}$ then $S$ is indeed right ample, a contradiction.
[…]
Theorem 3.5.
Let $B$ be a Brandt semigroup that is not a zero group. Then, for every $e\in E(B)\smallsetminus{0}$, the principle ideals $eB$ and $Be$ are, respectively, strict left and strict right ample subsemigroups of $B$.
Proof.
Let us arbitrarily fix an element $e=(\lambda,1,\lambda)\in E(B)\smallsetminus\{0\}$. We show that $eB$ is a strict left ample subsemigroup of $B$. (One may prove analogously that $Be$ is strict right ample in $B$.) Let $u$ be an arbitrary element of $eB$. If $u=0$, then $uu^{-1}=0\in eB$ and there is nothing to prove. So, assume that $u\in eB\smallsetminus\{0\}$. Then, we have
$u=(\lambda,1,\lambda)(j,h,l)=(\lambda,h,l),\,\text{ where }\,(j,h,l)\in B\,\text{ with }j=\lambda.$
Now, one can easily verify that $uu^{-1}=(\lambda,1,\lambda)$, where $u^{-1}$ denotes the inverse of $u$ in $B$. Clearly, $(\lambda,1,\lambda)$ is an element of $eB$. Hence, $eB$ is left ample (in $B$).
Next, we use Proposition 3.4 to prove that $S=eB$ is not right ample. The aim is to show that $\hat{\lambda}_{S}$ is not a monomorphism. Consider, for this purpose, $u=(\lambda,h,l)\in eB\smallsetminus\{0\}$. Let $u^{-1}$ denote the inverse of $u$ in some (up to isomorphism) inverse oversemigroup $T$ of $eB$, such that $u^{-1}u\in eB$ (if no such inverse semigroup exists then we are done). Then, after effectuating necessary identifications, one can easily verify from $uu^{-1}u=u\neq 0$ that $u^{-1}u=(l,1,l)$. Because $I$ is not a singleton, we can choose $u$ such that $l\neq\lambda$. But then,
$\begin{split}\hat{\lambda}_{u}&=(l,1,l)(eB)\longrightarrow(\lambda,h,l)(eB)\\ &=(l,1,l)(\lambda,1,\lambda)B\longrightarrow(\lambda,h,l)(\lambda,1,\lambda)B\\ &=0B\longrightarrow 0B\\ &=\hat{\lambda}_{0}\end{split},$
implying the $\hat{\lambda}_{S}$ is not a monomorphism. ∎
The following example provides a left ample semigroup of a Brandt semigroup that is not contained in any of the non-zero idempotent generated ideals.
Example 3.6.
Let $\mathbb{Q}^{+}$ be the multiplicative group of positive rationals, $\mathbb{N}_{1}$ the multiplicative semigroup of natural numbers starting from 1 and $I=\{1,2,3\}$. Consider the Brandt semigroup $B=(I\times\mathbb{Q}^{+}\times I)\cup\{0\}$ and its subsemigroup $S=(\{1,2\}\times\mathbb{N}_{1}\times I)\cup\{0\}$. Subsemigroup $S$ is left ample, because for any $n\in\mathbb{N}_{1}$, $j\in\{1,2\}$ and $i\in I$, we have $(j,n,i)(i,n^{-1},j)=(j,1,j)\in S$, but not right ample, because $(3,n^{-1},j)(j,n,3)=(3,1,3)\not\in S$. The non-zero idempotent generated ideals are $(i,1,i)B=(\{i\}\times\mathbb{Q}^{+}\times I)\cup\{0\}$. Clearly $S$ is not contained in any of them.
Remark 3.7.
The construction used in the above example works for any group (including finite ones). Let $G$ be a group and $I=\{1,2,3\}$. Consider the the Brandt semigroup $B=(I\times G\times I)\cup\{0\}$ and the subsemigroup $S=(\{1,2\}\times G\times I)\cup\{0\}$. (Actually we may take $S=(\{1,2\}\times H\times I)\cup\{0\}$, where $H$ is any submonoid of $G$.) Now we have, for any $g\in G$, $j\in\{1,2\}$ and $i\in I$, $(j,g,i)(i,g^{-1},j)=(j,1,j)\in S$, but $(3,g^{-1},j)(j,g,3)=(3,1,3)\not\in S$. Hence $S$ is strict left ample. On the other hand $S\not\subset(\{i\}\times G\times I)\cup\{0\}$ for any $i\in I$, but all principle ideals are of the form $(\{i\}\times G\times I)\cup\{0\}$. The reason for the previous example to work is that the strict ample subsemigroup (also a right ideal) is generated by two idempotents $(1,1,1)$ and $(2,1,2)$.
Problem 3.8.
Can we prove that Theorem 3.5 does not apply to the above example and remark?
Let $B=(I\times G\times I)\cup\{0\}$ be a Brandt semigroup. Then clearly every subsemigroup of $B$ with zero is of the form $S=(I^{\prime}\times H\times I^{\prime\prime})\cup\{0\}$ where $I^{\prime}$ and $I^{\prime\prime}$ are subsets of $I$ and $H$ is a subsemigroup of the group $G$; a subsemigroup without zero is of the form $I^{\prime}\times H\times I^{\prime\prime}$.
Theorem 3.9.
Let $S=(I^{\prime}\times G\times I^{\prime\prime})\cup\{0\}$ be a non-trivial subsemigroup of a Brandt semigroup $B=(I\times G\times I)\cup\{0\}$. Then $S$ is left (respectively, right) ample in $B$ if and only if $H$ is a left (respectively, right) ample submonoid of the group $G$.
Proof.
If $S$ is left ample in $B$ then for any non-zero element $s=(\lambda,a,\mu)\in S$ we have $ss^{-1}=(\lambda,a,\mu)(\mu,a^{-1},\lambda)=(\lambda,1,\lambda)\in S$. This implies that $aa^{-1}=1\in H$ for all $a\in H$, whence $H$ is a left ample submonoid of the group $G$.
Conversely if $H$ is a left ample submonoid of the group $G$ then for every $a\in H$ and its inverse $^{-1}\in G, aa$$a^{-1}\in H$. So $ss^{-1}=(\lambda,a,\mu)(\mu,a^{-1},\lambda)=(\lambda,aa^{-1},\lambda)\in S$, that is $S$ is left ample. A similar argument applies if ‘left ample’ is replaced by ‘right ample’. ∎
Theorem 3.10.
A subsemigroup $S=(I^{\prime}\times H\times I^{\prime\prime})\cup\{0\}$ of a Brandt semigroup $B=(I\times G\times I)\cup\{0\}$ is rich left (right) ample in $B$ if and only if $H$ is a subgroup of $G$.
Proof.
Suppose that $S$ is rich left ample in $B$. Take any $x=(i,a,\lambda),~y=(j,b,\mu)\in S$. Now $xy^{-1}\in S$ implies $(i,a,\lambda)(\mu,b^{-1},j)\in S$. Without loss of generality we may assume that $\lambda=\mu$ and thus $(i,ab^{-1},j)\in S$. This implies $ab^{-1}\in H$. Because $a$ and $b$ can be arbitrarily chosen, it follows that $H$ is a subgroup of $G$.
Conversely assume that $H$ is a subgroup of $G$. So $ab^{-1}\in H$ for every $a,b\in H$. Take any $x=(i,a,\lambda),~y=(j,b,\mu)\in S$. Then, $xy^{-1}$ is either $0\in S$ or equals $(i,ab^{-1},\mu)\in S$. Hence $S$ is rich left ample. The left version of the theorem can be proved similarly. ∎
4. Connection with dominions
Given a subsemigroup $S$ of semigroup $T$, an element $d\in T$ is said to be dominated by $S$ if $f|_{S}=g|_{S}$ implies $df=dg$ for all pairs of semigroup homomorphism $f,g:T\longrightarrow W$. The dominion of $S$ in $T$, denote by $Dom_{T}S$, is the set of all elements of $T$ dominated by $S$. In fact, $Dom_{T}S$ is a subsemigroup of $T$ and oversemigroup of $S$. Let $x\in Dom_{T}S$ for some inverse oversemigroup $T$ of $S$, and $W$ be a semigroup admitting homomorphisms $f,g:T\longrightarrow W$, such that $f|_{S}=g|_{S}$. Then, using the fact that $Imf$ and $Img$ are inverse subsemigroups of $W$, we have,
$(x^{-1})f=(xf)^{-1}=(xg)^{-1}=(x^{-1})g.$
This implies that $x^{-1}\in Dom_{T}S$. Hence $Dom_{T}S$ is an inverse subsemigroup of $T$.
Remark 4.1.
Let $S$ be a subsemigroup of an inverse semigroup $T$, and $\langle S\rangle_{T}$ denote the inverse subsemigroup of $T$ generated by $S$. Then $\langle S\rangle_{T}\subseteq Dom_{T}S$.
The following remark connects ampleness with dominions.
Remark 4.2 ([7], Proposition 1).
If $S$ be an ample subsemigroup of an inverse semigroup $T$ then $\langle S\rangle_{T}=Dom_{T}S$.
The next example demonstrates that the above condition is not sufficient.
Example 4.3.
Consider the inverse monoid $T=\mathbb{Q^{+}}$ of all positive rational numbers with usual multiplication. Then $S=\mathbb{N}\smallsetminus\{1\}$ is a non-ample subsemigroup of $T$; in fact, $T=\langle S\rangle_{T}$. Let $\alpha,\,\beta$ be homomorphisms from $T$ into a monoid $W$, such that $\alpha|_{S}=\beta|_{S}$. Then, one has, for all $n\in S$,
$\begin{split}\left(\frac{1}{n}\right)\alpha&=\left(2\cdot\frac{1}{2n}\right)\alpha=\left(2\right)\alpha\left(\frac{1}{2n}\right)\alpha=(2)\beta\left(\frac{1}{2n}\right)\alpha=\left(\frac{2n}{n}\right)\beta\left(\frac{1}{2n}\right)\alpha\\ &=\left(\frac{1}{n}\right)\beta(2n)\beta\left(\frac{1}{2n}\right)\alpha=\left(\frac{1}{n}\right)\beta(2n)\alpha\left(\frac{1}{2n}\right)\alpha=\left(\frac{1}{n}\right)\beta\left(\frac{2n}{2n}\right)\alpha\\ &=\left(\frac{1}{n}\right)\beta\left(1\right)\alpha=\left(\frac{1}{n}\right)\beta\left(1\right)\beta=\left(\frac{1}{n}\cdot 1\right)\beta=\left(\frac{1}{n}\right)\beta.\end{split}$
Because $\alpha$ and $\beta$ agree on its generators, it follows that they agree on $\langle S\rangle_{T}$. Hence $Dom_{T}S=\langle S\rangle_{T}$, as required.
Example 4.4.
Take $T=I_{X}$, where $X$ is an infinite set and let $S$ be its subsemigroup of full injective mappings on $X$. Clearly $T$ is non-group inverse semigroup and the subsemigroup $S$ is not ample in $T$. In fact $S$ is left ample but not right ample in $T$: For any $\in S, ss$$s^{-1}$ is the identity function on $X$ so is in $S$ but $s^{-1}s$ is not in $S$ as its domain is not necessarily whole of $X$. We show that $Dom_{T}(S)=S$ i.e., $S$ is closed in $T$ by using the zigzag argument. Take any $d\in Dom_{T}(S)\setminus U$ and let let $d=a_{0}y_{1}=x_{1}a_{1}y_{1}=x_{1}a_{2}y_{2}=x_{2}a_{3}y_{2}=x_{2}a_{4}$, where $a_{i}\in S,x_{i},y_{i}\in T$, be a zigzag in $T$ over $S$ with value $d$ (this is a zigzag of length 2 but the argument can be extended for a zigzag of any length). Since $a_{0}=x_{1}a_{1}$ so $X=dom(a_{0})=dom(x_{1})$ it follows that $x_{1}\in S$ and thus $x_{1}a_{1}=x_{2}a_{2}\in S$. Therefore $dom(x_{2}a_{2})=X$ which implies that $dom(x_{2})=X$ so $x_{2}\in S$. Hence $d=x_{2}a_{4}\in S$, showing that $S$ is closed in $T$. Thus $Dom_{T}(S)=S\subset\langle S\rangle_{T}.$
Acknowledgment. The authors are thankful to Professor Valdis Laan and dr. Ülo Reimaa for their valuable comments on the earlier drafts of this article. This research has been supported by Estonian Science Foundation’s grants PRG1204.